A deposit of $800 is placed in a savings account paying 6.9% annual interest compounded monthly. How much money will be in the account after 5 years?
Answer by stanbon(75887)

A deposit of $800 is placed in a savings account paying 6.9% annual interest compounded monthly. How much money will be in the account after 5 years?
-----------------------------
A(t) = P(1+(r/n))^(nt)
-----
A(5) = 800(1 + (0.069/12))^(12*5)
A(5) = 800(1.00575)^(60)
A(5) = 800(1.4106)
A(5) = $1128.48
====================
Cheers,
Stan H.